If the current one year interest rate is 3.5% and next year’s one year rate is...

  1. If the current one year interest rate is 3.5% and next year’s one year rate is 5.5%, what is the current two year rate?

  2. If the current three year interest rate is 7%, what is the current one year interest rate? Assume next year’s one year rate is 5.5% and the one year rate two years from now is 7.25%.

  3. Find the one year interest rate for four years from now. Assume the five year interest rate is 5.5%, the current one year rate is 3.5%, the one year rate one year from now is 5.75%, the one year rate two years from now is 4.5% and the one year rate three years from now is 6.75%.

  4. If a taxable bond has a yield to maturity of 5.5%, what is its after-tax yield?

    1. Tax rate = 20%

    2. Tax rate = 30%

    3. Tax rate = 40%

  5. Suppose you must choose between at taxable bond with a 7% yield and a municipal bond with a 4% yield. If your tax rate is 35%, which bond would you choose?

  6. If the municipal yield is 6%, what (pre-tax) taxable yield would a private bond issuer have to offer in order to make you indifferent between the two bonds if your tax rate is 45%?

  7. Given the following data, calculate the current five year interest rate: [Hint: There is more data here than you actually need.]

i1t

2.0%

i2t

2.8%

i1t+1

3.5%

i1t+2

3.0%

i1t+3

5.0%

i3t

2.8%

i1t+4

4.5%

i4t

3.4%

i1t+5

5.5%

I need help solving these with step by step instructions please!!
Answer #1


Question 1

The current two year rate would be the average of current one year interest rate and next year's one year rate.

Calculate the current two year rate -:

Current two year rate = [Current one year interest rate + Next year's one year interest rate]/2

Current two year rate = [3.5 + 5.5]/2 = 9/2 = 4.5%

Thus,

The current two year rate is 4.5%